please help me simplify the following expression:
a+1/(a^2-2a-3)- a+1/(a^2+4a+3)
i have tried to factorise the denominators to find common factors but cannot come up with the answer in my book: 6/(a-3)(a+3). Also, tried to put the expression into the automatic calculator on here but it came out with a different answer. Help! Thanks, Ade

Answer by Alan3354(69443)

a+1/(a^2-2a-3)- a+1/(a^2+4a+3)
----------------
= a+1/(a-3)*(a+1) - a+1/(a+1)*(a+3)
= 1/(a-3) - 1/(a+3)
= ((a+3) - (a-3))/(a^2 - 9)
= 6/(a^2 - 9)